OpenStudy (anonymous):

486 x 736 x 73-854=3

OpenStudy (aum):

"How is the graph of the function y - (x - 3)^2" Is that supposed to be y = (x - 3)^2 (= instead of -) ?

OpenStudy (aum):

Replacing "x" by "x-a" in a function shifts the function to the RIGHT by "a" units. Replacing "x" by "x+a" in a function shifts the function to the LEFT by "a" units.

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

Thank you.

OpenStudy (aum):

You are welcome.